OpenStudy (tukitw):

A hot-air balloonist drops an apple over the side while the balloon is accelerating upward at 4.0 ms^-2 during lift-off. If the speed of the balloon is 2.0 ms^-1 at the moment of release, how long does it take for the apple to reach 200 m below the balloon?

OpenStudy (shamim):

h= ut+1/2 gt ^2

OpenStudy (shamim):

h=200 u=-2 g=9.8

OpenStudy (tukitw):

Do I need to consider the displacement of the balloon? Since its going up while the apple's falling.

OpenStudy (shamim):

I think the velocity of apple will b equal to the velocity of baloon

OpenStudy (shamim):

so the velocity of apple will b 2 ms^-1

OpenStudy (shamim):

and it will b upward

OpenStudy (shamim):

ya u can consider the displacement of baloon

OpenStudy (tukitw):

So it's displacement of balloon + displacement of apple = 200?

OpenStudy (shamim):

baloon will go upward but apple will go first upward and then it will go downward

OpenStudy (shamim):

actually the apple will go more than 200 m

OpenStudy (tukitw):

Displacement of balloon - Upwards displacement of apple + Downwards displacement of apple = 200?

OpenStudy (shamim):

the displacement of apple will b 200 m

OpenStudy (shamim):

ya right

OpenStudy (tukitw):

Like this? -2t - 2t^2 - (-2t + 4.9t^2) = 200

OpenStudy (shamim):

no

OpenStudy (shamim):

200=-2 t+4.9 t^2

OpenStudy (shamim):

h= ut+(1/2) gt ^2

OpenStudy (tukitw):

So t = 6.60 s?

OpenStudy (shamim):

may b

OpenStudy (shamim):

I did not calculate

OpenStudy (tukitw):

Thanks

OpenStudy (shamim):

welcome